A rocket launched vertically upward sort up 60kg of exhaust gasses every second with a velocity of 200ms 1. What is the upward force on the rocket. 2.if the rocket has the Mass of 500kg, what is the initial acceleration upward
1
Expert's answer
2019-05-24T09:56:38-0400

1.


"F=v\frac{dm}{dt}=200\frac{60}{1}=12000\ N=12\ kN"

2.


"a=\frac{F}{M}=\frac{12000}{500}=24\frac{m}{s^2}"